### v1

# Large-Scale Text-to-Image Model with Inpainting is a Zero-Shot Subject-Driven Image Generator

Abstract
* Correspondence to: Sungroh Yoon (sryoon@snu.ac.kr)
Subject-driven text-to-image generation aims to produce images of a new subject within a desired context by accurately capturing both the visual characteristics of the subject and the semantic content of a text prompt. Traditional methods rely on time- and resource-intensive fine-tuning for subject alignment, while recent zero-shot approaches leverage on-the-fly image prompting, often sacrificing subject alignment. In this paper, we introduce Diptych Prompting, a novel zero-shot approach that reinterprets as an inpainting task with precise subject alignment by leveraging the emergent property of diptych generation in large-scale text-to-image models. Diptych Prompting arranges an incomplete diptych with the reference image in the left panel, and performs text-conditioned inpainting on the right panel. We further prevent unwanted content leakage by removing the background in the reference image and improve fine-grained details in the generated subject by enhancing attention weights between the panels during inpainting. Experimental results confirm that our approach significantly outperforms zero-shot image prompting methods, resulting in images that are visually preferred by users. Additionally, our method supports not only subject-driven generation but also stylized image generation and subject-driven image editing, demonstrating versatility across diverse image generation applications.
Figure 1: Given a single reference image, our Diptych Prompting performs zero-shot subject-driven text-to-image generation through diptych inpainting. Building on the (a) diptych generation capability of FLUX [21], we extend it to diptych inpainting with a separate module, resulting in (b) versatility across various tasks including subject-driven text-to-image generation, stylized image generation, and subject-driven image editing.
1 Introduction
With recent advancements in generative models, text-to-image (TTI) models [36, 33, 3, 8, 6, 40, 2, 9] have significantly improved, enabling the generation of photorealistic images based on text prompts. Beyond generating images from text, these models support various text-based image tasks, including text-guided editing [12, 28, 16, 48, 11], text-guided style transfer [13, 38, 41], and subject-driven text-to-image generation [39, 20, 10, 50, 22, 47, 30, 46, 26, 45, 31]. Specifically, subject-driven text-to-image generation aims to synthesize images of a specific subject in various contexts based on a text prompt and a reference image, while achieving both subject and text alignment.
Early research on subject-driven text-to-image generation [10, 39, 20, 45] enables the model to synthesize a new subject through fine-tuning on a small set of images containing the target subject. While they achieve strong subject alignment via optimization, they are time- and resource-intensive, requiring hundreds of iterative steps of optimization for each new subject. As an alternative, zero-shot approaches [50, 22, 47, 30, 46, 26, 31] have emerged that do not require additional fine-tuning and instead utilize image prompting through a specialized image encoder. These methods extract the image feature from a reference image and integrate it into the TTI model alongside the text feature. While they achieve on-the-fly subject-driven text-to-image generation with a single forward pass of the encoder, these encoder-based image prompting frameworks suffer from unsatisfactory subject alignment, particularly in capturing granular details.
Recently, as models in NLP fields have been scaled up and demonstrated remarkable capabilities [4, 1], large-scale TTI models [8, 21] have similarly emerged. Notably, the recently released model, FLUX [21], has demonstrated exceptional text comprehension and the ability to effectively translate this understanding into images, even for highly complex and lengthy texts. Among the various capabilities of FLUX, we focus on its ability to generate high-quality diptychs--two-paneled art pieces in which each panel contains an interrelated image. As shown in Fig. 1 (a), FLUX’s advanced text understanding and high-resolution image generation enable it to generate side-by-side images of the same object, each reflecting a different context as specified in the prompt for each panel.
Motivated by FLUX’s ability to generate diptychs, we propose “Diptych Prompting”, a novel inpainting-based framework for zero-shot, subject-driven text-to-image generation. In our approach, we reinterpret the task as a diptych inpainting process: the left panel contains a reference image of the subject as a visual cue, and the right panel is generated through inpainting based on a text prompt describing the diptych with the desired context. Using text-conditioned diptych inpainting, Diptych Prompting aligns the generated image in the right panel with both the reference subject and the text prompt. We enhance this process by removing the background from the reference image to prevent content leakage and focus solely on the subject, and by enhancing attention weights between panels to ensure fine-grained details preservation. These two components enable Diptych Prompting to achieve more consistent, high-quality subject-driven text-to-image generation.
Through various experiments, Diptych Prompting demonstrates superior performance over existing encoder-based image prompting methods, more effectively capturing both subject and text and producing results preferred by human evaluators. Additionally, our method is not limited to subjects; it can also be applied to styles, enabling stylized image generation [41, 38, 13] when a personal style image is provided as a reference. Furthermore, we showcase the extensibility of our approach to subject-driven image editing [49], allowing modification of specific regions in the target image with the reference subject. By arranging the target image in the right panel of diptych and masking only the region for editing in Diptych Prompting, we successfully integrate the reference subject into the target image.
Our contributions can be summarized as follows:
  • To the best of our knowledge, we are the first to approach zero-shot subject-driven text-to-image generation as an inpainting task, offering a new perspective by highlighting the diptych generation capabilities of FLUX.
  • We propose two techniques to prevent content leakage and reliably capture details in the target subject: isolating the subject from its background and enhancing attention weights between panels.
  • We validate our method’s versatility and robustness, extending its effectiveness even to style-driven generation and subject-driven image editing through comprehensive qualitative and quantitative results.
2 Related Works
2.1 Diffusion-based Text-to-Image Models
Diffusion models [15, 42, 43, 18] have led to significant advancements in TTI models, including GLIDE [29], LDM [36], DALL-E 2 [35], Imagen [40], and eDiff-I [2]. Among these, the Stable Diffusion (SD) series [36, 33, 8] has gained particular attention for its open-source nature and competitive performance to previous research. Starting with the v1 model, which utilizes a U-Net [37] architecture with cross-attention for text, it evolved through v2 and then to SD-XL [33], with improvements in dataset scale, model architecture, resolution, and generation quality.
Recently, generative model research [32] has achieved notable performance improvement by incorporating transformer [44] architectures into diffusion models instead of U-Net. Driven by this advancement, emerging studies now integrate transformer architecture into TTI models, most notably SD-3 [8] and FLUX [21]. Both models employ the MultiModal-Diffusion Transformer (MM-DiT) architecture, an advanced design for TTI models that conducts joint attention on concatenated text and image embeddings,
$Q=[Q_{t};Q_{i}],K=[K_{t};K_{i}],V=[V_{t};V_{i}],$ (1)
$\text{A}(Q,K,V)=W(Q,K)V=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V,$ (2)
where [;][;][ ; ] is the concatenation, $Q$, $K$, and $V$ represent the key components of attention--query, key, and value, respectively; $W$ is the attention weight, and $A$ is the output of the attention. FLUX, in particular, is the largest-scale TTI model among open-source models and exhibits advanced performance in both text comprehension and image generation quality, surpassing previous open-source models.
2.2 Text-Conditioned Inpainting
Image inpainting aims to fill the missing regions of an incomplete image $I$ using a binary mask $M$ that specifies the areas to be reconstructed. Recent advancements in TTI models have led to the development of text-conditioned inpainting [49], which completes the missing regions to align not only with the visible region but also with a text prompt,
$\hat{I}=F_{\theta}(I,M,T),$ (3)
where $T$ is the text describing the desired context, and $F_{\theta}$ is the generation process of the text-conditioned models. Various methods [43, 49] have been proposed to implement a plausible $F_{\theta}$ from the pre-trained TTI models.
While an early approach [43] employs pre-trained diffusion models without any further training, more recent works fine-tune the pre-trained TTI model or train additional modules [51] specifically for inpainting tasks. Through additional training for inpainting, these models achieve the two main objectives of text-conditioned inpainting: alignment with the visible regions in $I$ and alignment with the text prompt. Among various inpainting modules, ControlNet [51] equips FLUX with inpainting capability, providing inpainting-specific conditioning for enhanced control. By leveraging this module, we interpret inpainting as a framework for subject-driven text-to-image generation.
2.3 Subject-Driven Image Generation
There has been extensive research on subject-driven text-to-image generation [39, 20, 10, 50, 22, 47, 30, 46, 26, 45, 31], where the generated images not only render the various contexts described by the text prompt but also include the specific subject according to reference images. Subject-driven text-to-image generation is generally categorized into two groups based on whether they require additional training for each new subject.
The first category [39, 20, 10, 45] involves fine-tuning on a small set of subject images (e.g., $3$-$5$ images) to learn the visual subject and how to generate it. While these methods achieve strong subject alignment through optimization on the subject, the fine-tuning requires retraining for each new subject, making them time- and resource-intensive. Moreover, optimizing on a small set of images may lead to overfitting on the new subject and catastrophic forgetting of prior knowledge which should be carefully prevented.
The second group [50, 22, 47, 30, 46, 26, 31] addresses the limitations of the first by introducing image prompting, a framework that incorporates a reference image to guide the generated output. In this approach, a specialized image enco der [34] is utilized to support an additional image input alongside the text prompt, allowing the diffusion model to capture the characteristics of the reference subject while adhering to the context described by the text prompt. This approach enables zero-shot subject-driven text-to-image generation; however, it often lacks fidelity in the synthesized subject. We propose a novel inpainting-based zero-shot approach for subject-driven text-to-image generation using a large-scale TTI model with an inpainting module.
3 Method
Figure 2: Diptych Generation Comparisons. We generate the diptych images with various TTI models from the following diptych text: “A diptych with two side-by-side images of same cat. On the left, a photo of a cat in front of Eiffel Tower. On the right, replicate this cat exactly but as a photo of a cat in the jungle”.
Figure 3: (a) Overall Diptych Prompting Framework. Given the incomplete diptych $I_{\text{diptych}}$, text prompt $T_{\text{diptych}}$ describing the diptych, and the binary mask $M_{\text{diptych}}$ specifying the right panel as the inpainting target, FLUX with ControlNet module performs text-conditioned inpainting on the right panel while referencing the subject in the left panel. (b) Reference Attention Enhancement. To capture the granular details of the subject in left panel, we enhance the reference attention, an attention weight between the query of the right panel and the key of the left panel.
3.1 Diptych Generation of FLUX
A ‘Diptych’ is an art term referring to a two-paneled artwork in which two panels are displayed side by side, each containing interrelated content. Previous work, HQ-Edit [17], proposed a pipeline for creating an image editing dataset in the form of diptychs using DALL-E 3 [3]. The strong text-image alignment of the large-scale TTI model, DALL-E 3, plays a critical role in creating coherent editing pairs in the diptych.
The recently released large-scale open-source TTI model, FLUX [21], demonstrates strong text comprehension and image generation capabilities, even surpassing DALL-E 3 [3]. Notably, its capabilities also extend to diptych generation: when we generate an image with diptych text $T_{\text{diptych}}$, “A diptych with two side-by-side images of the same {object}. On the left, {description of left image}. On the right, replicate this {object} but as {description of right image}, FLUX synthesizes a diptych image where the subjects in each panel are interrelated and each description of panel is accurately represented, as shown in Fig. 2.
Generating high-quality diptych images requires the robust text-image alignment capability of large-scale TTI model, in which smaller models fall short. Compared to previous models such as SD-v2 [36], SD-XL [33], and SD-3 [8], only FLUX [21] successfully synthesizes accurate diptych images that not only effectively interrelate subjects across panels but also render the correct contexts for each panel described in the diptych text. Therefore, we choose FLUX as the base model for our proposed methodology due to its superior ability to generate accurate and contextually aligned diptych images.
3.2 Diptych Prompting Framework
For zero-shot subject-driven text-to-image generation, most approaches rely on a specialized image encoder for image prompting that extracts image feature from a reference image and integrates it into the TTI model. Instead, to inject detailed subject characteristics into the generated image in a zero-shot manner, we propose a novel prompting approach that reinterprets zero-shot method from the perspective of inpainting, as illustrated in Fig. 3 (a).
Given the reference subject image and the target text prompt describing the desired context, Diptych Prompting begins with the triplets for inpainting-based prompting: an incomplete diptych image $I_{\text{diptych}}$, a binary mask $M_{\text{diptych}}$ specifying the missing region, and a diptych text $T_{\text{diptych}}$.
For the incomplete diptych image $I_{\text{diptych}}$, we concatenate two images along the width dimension with the left panel containing the reference subject image and the right panel consisting of a blank image of the same size to be inpainted. We observe that simple diptych inpainting often results in excessive interrelation with the reference image by mirroring even subject-unrelated contents, such as background, pose, and location (Fig. 4). To prevent this, we remove the background of the reference image through the background removal process $G_{\text{seg}}$ using Grounding DINO [25] and Segment Anything Model (SAM) [19]. In this process, Grounding DINO uses the subject name to acquire a bounding box of target subject through grounded object detection, and SAM performs subject segmentation with this detection box and removes the background, preparing it as the left panel,
$I_{\text{diptych}}=[G_{\text{seg}}(I_{\text{ref}});~{}\emptyset~{}].$ (4)
Figure 4: Background Removal Effects. Simple diptych inpainting exhibits content leakage from the reference image, including background, pose, and location. We mitigate this unwanted leakage through background removal by $G_{\text{seg}}$.
Additionally, our binary mask $M_{\text{diptych}}$ designates the location of the reference image in the left panel with zeros to provide visual cues, while marking the right panel with ones to indicate the missing areas to be filled,
$M_{\text{diptych}}=[\mathbf{0}_{h\times w};\mathbf{1}_{h\times w}],$ (5)
where $\mathbf{0}_{h\times w}$ and $\mathbf{1}_{h\times w}$ have the same size with each corresponding panel.
For the diptych text $T_{\text{diptych}}$ describing the diptych configuration with desired context, we utilize the prompt template used in Sec. 3.1. From the target text prompt, we use the subject name of reference subject for object, resulting in the following final diptych text: “A diptych with two side-by-side images of same {subject name}. On the left, a photo of {subject name}. On the right, replicate this {subject name} exactly but as {target text prompt}.
Using these triplets, our Diptych Prompting performs the text-conditioned inpainting,
$\hat{I}_{\text{diptych}}=[G_{\text{seg}}(I_{\text{ref}});\hat{I}_{\text{gen}}]=F_{\theta}(I_{\text{diptych}},M_{\text{diptych}},T_{\text{diptych}}),$ (6)
where $\hat{I}_{\text{gen}}$ represents the desired subject-driven image.
3.3 Reference Attention Enhancement
Diptych Prompting reconstructs the right panel of the diptych by referencing the subject in the left panel. However, FLUX [21] with inpainting module often struggles to fully capture the fine details of the subject.
Recent studies [12, 13, 7] have shown that the image generation process in U-Net-based TTI models can be controlled by manipulating key components of the attention--query, key, value, and attention weight--yet similar techniques remain largely unexplored in transformer-based architectures. Given that FLUX, built on the MM-DiT architecture, incorporates more attention blocks than previous U-Net-based models, it offers greater potential for such control. In Diptych Prompting, we note that FLUX synthesizes both the reference and generated image simultaneously in a diptych format through its attention blocks and computes the attention between the left and right panels. This leads us to enhance reference attention--the influence of the left panel on the right--to better capture granular details of the reference subject.
In the attention blocks of FLUX, the image feature part can be divided into two regions in diptych inpainting, corresponding to the left and right panel,
$Q=[Q_{t};Q_{li};Q_{ri}],K=[K_{t};K_{li};K_{ri}],$ (7)
where $(\cdot)_{t}$ is the feature for text, $(\cdot)_{li}$ is for the left panel, and $(\cdot)_{ri}$ is for the right panel.
From this, we acquire the attention weight $W(Q,K)$ as shown in Eq. 2, where $W(Q,K)\in\mathbb{R}^{(l_{t}+l_{li}+l_{ri})\times(l_{t}+l_{li}+l_{ri})}$, $l_{t}$ is the text sequence length, and $l_{\cdot i}$ is sequence length of each panel in attention blocks as described in Fig. 3 (b). We enhance the reference attention, the attention weight between the query of right panel ($Q_{ri}$) and the key of left panel ($K_{li}$) by rescaling the submatrix $W(Q_{ri},K_{li})$ with $\lambda>1$.
4 Experiments
4.1 Experimental Settings
Implementation Details Our method is implemented based on the large-scale TTI model, FLUX-dev [^1] FLUX.1-dev: https://huggingface.co/black-forest-labs/FLUX.1-dev, with the additional ControlNet-Inpainting module222 FLUX.1-dev-Controlnet-Inpainting-Beta: https://huggingface.co/alimama-creative/FLUX.1-dev-Controlnet-Inpainting-Beta. We perform diptych inpainting on a canvas with an aspect ratio of $1$:$2$, sized at $768\times 1536$, where the left half ($768\times 768$) serves as the reference. During inference, the ControlNet conditioning scale is set to $0.95$ and the reference attention rescaling parameter $\lambda$ is set to $1.3$ for diptych inpainting performed over $30$ steps and a guidance scale of $3.5$ [14, 27].
Evaluations We measure zero-shot subject-driven text-to-image generation performance on DreamBench [39] that contains 30 subjects, each with 25 evaluation prompts. Following previous work [39], we generate 4 images per subject and prompt, resulting in a total of $3000$ images. These images are evaluated using the DINO [5] and CLIP [34]-based metrics which quantify the two objectives of subject-driven text-to-image generation: subject alignment and text alignment. Subject alignment is measured by the average pairwise cosine similarity of features between generated images and real images using the DINO and CLIP image encoders (DINO, CLIP-I). Text alignment is measured by the pairwise cosine similarity between the CLIP image embeddings of the generated images and the CLIP text embeddings of the target texts (CLIP-T).
Figure 5: Qualitative Comparisons. Please zoom in for a more detailed view and better comparison.
4.2 Baseline Comparisons
We compare our method to previous zero-shot subject-driven text-to-image methods with encoder-based image prompting, including ELITE [47], BLIP-Diffusion [22], Kosmos-G [30], Subject-Diffusion [26], IP-Adapter [50], MS-Diffusion [46], and $\lambda$-Eclipse [31]. The details of these models are provided in Appendix A.
Qualitative Results Our qualitative results are presented in Fig. 5, where the reference images are at the leftmost column and generation results are at the right. Despite using an inpainting approach without any specialized training for subject-driven text-to-image generation, Diptych Prompting generates high-quality samples and accurate renderings of text prompt across diverse subjects and situations, significantly outperforming results compared to previous approaches. Our method also demonstrates impressive performance in capturing the granular details of reference subjects, even with challenging examples containing characteristic fine details, such as a ‘monster toy’ or ‘backpack’.
Human Preference Study We confirm the outstanding performance of our method in terms of human perception through a human preference study. We conduct a paired comparisons of our method with each baseline from two perspectives: subject alignment and text alignment. Using Amazon Mechanical Turk, we collected $450$ responses from $150$ participants for each baseline and each perspective. As shown in Tab. 1, Diptych Prompting outperforms all baselines by a large margin ($p<0.01$ in the Wilcoxon signed-rank test), which is consistent with the qualitative results. Detailed information and full instructions about our human preference study are included in Appendix C.
Subject Align (%) Text Align (%)
Method win tie lose win tie lose
ELITE [47] 77.9 4.3 17.8 75.2 8.6 16.2
BLIP-Diff [22] 73.8 8.6 17.6 77.8 4.3 17.9
$\lambda$-Eclipse [31] 80.4 4.2 15.4 74.0 3.3 22.7
MS-Diff [46] 59.3 15.6 25.1 58.9 9.1 32.0
IP-A (SD-XL) [50] 76.2 9.7 14.1 76.2 9.7 14.1
IP-A (FLUX) [50] 69.8 12.0 18.2 65.2 20.6 14.2
Table 1: Human Preference Study. We report results of pairwise comparisons between Diptych Prompting and publicly available baselines in two aspects: subject alignment and text alignment. ‘IP-A’ denotes the abbreviation for IP-Adapter.
Quantitative Results For quantitative aspects, the comparison results are in Tab. 2. Diptych Prompting demonstrates comparable or superior performances in both subject alignment and text alignment, as measured by DINO and CLIP-T scores. We also note that all baseline methods perform image prompting using the CLIP image encoder, resulting in high CLIP-I scores. In contrast, our inpainting-based zero-shot approach leverages the inherent generation capability of a large-scale TTI model without specialized image encoder, which presents a slight disadvantage in terms of CLIP-I. However, the results from other metrics, qualitative comparisons, and human evaluation studies across both aspects confirm the effective performance and robustness of our method.
Method Model DINO CLIP-I CLIP-T
ELITE [47] SD-v1.4 0.621 0.771 0.293
BLIP-Diff [22] SD-v1.5 0.594 0.779 0.300
Kosmos-G [30] SD-v1.5 0.694 0.847 0.287
Subject-Diff [26] - 0.711 0.787 0.303
$\lambda$-Eclipse [31] Kan-v2.2 0.613 0.783 0.307
MS-Diff [46] SD-XL 0.671 0.792 0.321
IP-Adapter [50] $\dagger$ SD-XL 0.613 0.810 0.292
IP-Adapter [50$\ddagger$ FLUX 0.561 0.725 0.351
Diptych Prompting FLUX 0.688 0.758 0.345
Table 2: Quantitative Comparisons. We compare our method to encoder-based image prompting methods in three metrics. $\dagger$ denotes the obtained value from [31], and $\ddagger$ indicates our re-evaluation with publicly available weights.
4.3 Ablation Studies
Model Inpainting Scale DINO CLIP-I CLIP-T
SD-3 Zero-shot - 0.475 0.670 0.330
ControlNet 0.95 0.576 0.699 0.326
FLUX Zero-shot - 0.555 0.720 0.336
ControlNet 0.5 0.628 0.737 0.351
0.8 0.670 0.750 0.349
0.95 0.689 0.758 0.344
Table 3: Model Selection. We present an ablation results of various base models, inpainting method, and the ControlNet conditioning scale for Diptych Prompting.
$G_{seg}$ $\lambda$ DINO CLIP-I CLIP-T
1.3 0.759 0.783 0.333
1.0 0.647 0.745 0.343
1.3 0.688 0.758 0.345
1.5 0.670 0.750 0.342
Table 4: $\bm{G_{\text{seg}}}$ and $\bm{\lambda}$ Ablation. We report the ablation results of background removal and reference attention enhancement.
To analyze the factors contributing to the performance, we conduct in-depth ablation studies for Diptych Prompting.
Model Selection We validate our method across various base models and inpainting methods including the zero-shot approach [43]. As shown in Tab. 3, we demonstrate that utilizing a high-capacity base model and enhancing the inpainting method leads to improved zero-shot subject-driven text-to-image generation. From these results, we employ the combination of a robust base model, an effective inpainting method, and an appropriate inpainting-conditioning scale for Diptych Prompting. Integrating advanced base models or inpainting methods is expected to improve the performance and expand our method to more tasks in the future.
$\bm{G_{\text{seg}}}$ and $\bm{\lambda}$ Ablation We conduct additional ablation experiments to verify the effectiveness of background removal in preventing the content leakage and reference attention enhancement in the fine-grained details preservation in Diptych Prompting, as shown in Tab. 4.
When background removal is not applied, we observe copy-and-paste-like results (Fig. 4). These results cause subject alignment metrics to increase significantly due to the mirroring of the reference image, yet at the expense of text alignment, resulting in higher DINO, CLIP-I scores and lower CLIP-T score.
We also assess the impact of varying the rescaling factor $\lambda$ on reference attention enhancement. Rescaling attention weights between the right panel query and the left panel key helps to capture fine details, thereby improving subject alignment. However, using too high values introduces excessive inductive bias, causing abnormal attention weights that negatively impact performance. Qualitative transitions with respect to $\lambda$ can be verified in Appendix F.
4.4 Applications
Figure 6: Qualitative Comparisons of Stylized Image Generation. Using a style image as a reference, Diptych Prompting generates stylized images.
With the strong capabilities demonstrated by Diptych Prompting, we also explore how it can be applied to tasks beyond subject-driven text-to-image generation.
Stylized Image Generation We extend our method beyond subject images and perform stylized image generation using various style images as references. Using style images and prompts in StyleDrop [41], we employ Diptych Prompting as same, but replace the subject name with the term ‘style’ in diptych text and without attention enhancement ($\lambda=1$) for referencing only the stylistic elements except the content. Diptych Prompting successfully generates the stylistic image reflecting the style of the reference as shown in Fig. 6, and the quantitative comparisons are available in Appendix G.
Subject-Driven Image Editing We further adapt our approach to support inpainting-based subject-driven image editing that modifies the target image with the specific subject. In this setup, we utilize Diptych Prompting with reference subject image, yet assign the right panel as the editing target image and apply the mask only to the region to be edited. Editing results are shown in Fig. 7. Owing to the capability of Diptych Prompting, edited images effectively preserve unmasked areas while seamlessly integrating the desired subject within the target region.
Figure 7: Subject-Driven Image Editing. Diptych Prompting extends to subject-driven image editing by placing the target image on the right panel and masking only the area to be edited.
5 Conclusion
In this paper, we proposed Diptych Prompting, an inpainting-based approach for zero-shot subject-driven text-to-image generation. Diptych Prompting performed text-conditioned diptych inpainting: the left panel is a reference image containing the subject, and the right panel is inpainted based on a text prompt that describes the diptych containing the desired context. By removing the background and enhancing reference attention, we eliminated unnecessary content leakage and improved subject alignment. This innovative approach enjoyed the inherent properties of large-scale TTI models, achieving superior results over previous methods, particularly in accurately capturing target subjects and representing complex contexts. We also demonstrated the versatility of our method in stylized image generation and subject-driven image editing. Building on these contributions, we anticipate that Diptych Prompting will inspire new directions in image generation and across a wide range of generative tasks, including video and 3D.
Acknowledgement
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) [No. 2022R1A3B1077720], Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)], and the BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University in 2024.
Appendix
Appendix A Baselines
We provide the details of the encoder-based image prompting baselines that we compared in human preference study, as well as in qualitative and quantitative evaluations. All of them utilize a specialized image encoder which extracts image feature from the reference image and injects it into the TTI model. While these models train the specialized image encoder to enable image prompting for zero-shot subject-driven text-to-image generation, they compromise subject alignment, especially in the granular details of the subject. For qualitative results and the human preference study, we compare our method only to the baselines with available open-source weights.
  • ELITE [^3] ELITE: https://github.com/csyxwei/ELITE [47] encodes the visual concepts into textual embeddings, leveraging global and local mapping networks to represent primary and auxiliary features separately, ensuring high fidelity and editability in subject-driven text-to-image generation.
  • BLIP-Diffusion [^4] BLIP-Diff: https://github.com/salesforce/LAVIS/tree/main/projects/blip-diffusion [22] pre-trains a multimodal encoder following BLIP-2 [23] which produces the text-aligned visual representation of the target subject, and learns the subject representation to enable the TTI model to perform efficient subject-driven text-to-image generation.
  • Kosmos-G [30] aligns the output space of Multimodal Large Language Models (MLLMs) with the CLIP [34] space by anchoring the text modality, and bridges the MLLM with a frozen TTI model using AlignerNet and instruction tuning. As there are no available weights for this baseline, we cannot conduct the human preference study and can only compare using automatic quantitative metrics based on the values reported in their paper.
  • Subject-Diffusion [26] utilizes an image encoder trained on their own large-scale subject-driven dataset to incorporate both coarse and fine-grained reference information into the pre-trained TTI model, enabling high-fidelity subject-driven text-to-image generation without test-time fine-tuning. Subject-Diffusion also has no available open-source weights, so we only conduct the quantitative comparisons with their reported values in the paper.
[…]
  • MS-Diffusion [^6] MS-Diff: https://github.com/MS-Diffusion/MS-Diffusion [46] introduces a layout-guided framework for multi-subject zero-shot subject-driven text-to-image generation by employing a grounding resampler for detailed feature integration and a multi-subject cross-attention mechanism to ensure spatial control and mitigate subject conflicts.
  • IP-Adapter [^7] IP-Adapter (SD-XL): https://huggingface.co/h94/IP-Adapter  [^8] IP-Adapter (FLUX): https://huggingface.co/XLabs-AI/flux-ip-adapter [50] trains an effective lightweight adapter to enable image prompting for pre-trained TTI models, using a decoupled cross-attention mechanism with separate cross-attention layers for text and image prompts. At the time the IP-Adapter paper was released, SD-v1.5 [36] was used; however, more recent versions, including SD-XL [33], SD-3 [8], and FLUX [21], have since been made available. For quantitative comparisons, we referenced the results for the SD-XL version from another study [31], while we conducted our own evaluations for the FLUX version to ensure a fair comparison. In all experiments using IP-Adapter, regardless of the base model version, the conditioning scale is set to $0.6$.
Figure A1: DreamBooth Comparisons. Quantitative comparisons to DreamBooth-LoRA with various rank values.
Appendix B Subject-Driven Text-to-Image Generation
B.1 Evaluation Setting
We conduct the main comparisons with baselines on $30$ subjects in DreamBench [39]. These consist of $21$ objects and $9$ live subjects, with $25$ evaluation prompts for the objects or live subjects. Diptych Prompting uses the subject name to refer to the target subject and utilizes evaluation prompts that include the subject name for the target description in diptych text. In all zero-shot baselines and our method, we enhance the subject names by adding descriptive modifiers to more accurately refer to the target subjects in the text prompt. The subject names for each subject are summarized as follows in the form of (directory name, subject name):
  • backpack, backpack
  • backpack_dog, backpack
  • bear_plushie, bear plushie
  • berry_bowl, ‘Bon appetit’ bowl
  • can, ‘Transatlantic IPA’ can
  • candle, jar candle
  • cat, tabby cat
  • cat2, grey cat
  • clock, number ‘3’ clock
  • colorful_sneaker, colorful sneaker
  • dog1, fluffy dog
  • dog2, fluffy dog
  • dog3, curly-haired dog
  • dog5, long-haired dog
  • dog6, puppy
  • dog7, dog
  • dog8, dog
  • duck_toy, duck toy
  • fancy_boot, fringed cream boot
  • grey_sloth_plushie, grey sloth plushie
  • monster_toy, monster toy
  • pink_sunglasses, sunglasses
  • poop_emoji, toy
  • rc_car, toy
  • red_cartoon, cartoon character
  • robot_toy, robot toy
  • shiny_sneaker, sneaker
  • teapot, clay teapot
  • vase, tall vase
  • wolf_plushie, wolf plushie
B.2 Comparison with Fine-Tuning-Based Method
To provide a more comprehensive comparison, we also compare with DreamBooth [39], a representative fine-tuning-based method. For efficient training, we attach a LoRA adapter to the pre-trained FLUX and perform fine-tuning by training only the LoRA adapter while freezing the FLUX. We train for 300 steps using the Adam optimizer with a learning rate of $1\times 10^{-4}$. Additionally, to compare different fine-tuning model capacities, we adjusted the rank of the LoRA adapter and conducted comparative experiments using the same metrics (DINO, CLIP-I, CLIP-T). The results are presented in Fig. A1, where our Diptych Prompting demonstrates superior performance across various model capacities.
B.3 Additional Results
We include additional samples of Diptych Prompting in Fig. A2 and Fig. A3 for diverse objects and contexts. As demonstrated in the results, our methodology achieves high-quality image generation and satisfies both subject alignment and text alignment in a zero-shot manner by leveraging FLUX’s capabilities. Notably, this is accomplished without any specialized training for subject-driven text-to-image generation. We also note that the fine details in the target subject are well reflected in the generated results, even for challenging subjects that previous zero-shot methods struggled with (e.g., robot toy, ‘Bon appetit’ bowl).
Appendix C Human Preference Study
Following the previous work [39], we perform the human preference study by pairwise comparison in two separate questionnaires for each aspect: subject alignment and text alignment. In both questionnaires, users are presented with a reference image, a target text, and two images generated by each method. They are then asked to select which image better satisfies the desired objective according to the following instructions.
For subject alignment:
  • Inspect the reference subject and then inspect the generated subjects.
  • Select which of the two generated items reproduces the identity (item type and details) of the reference item
  • The subject might be wearing accessories (e.g., hats, outfits). These should not affect your answer. Do not take them into account.
  • If you’re not sure, select Cannot Determine / Both Equally.
  • Which Machine-Generated Image best matches the subject of the reference image?
For text alignment:
  • Inspect the target text and then inspect the generated items.
  • Select which of the two generated items is best described by the target text.
  • If you’re again not sure, select Cannot Determine / Both Equally.
  • Which Machine-Generated Image is best described by the reference text?
Appendix D Diptych Generation
Model Arch Param DINO CLIP-I CLIP-T
SD-v2 U-Net 1.2B 0.504 0.744 0.260
SD-XL U-Net 3.5B 0.941 0.954 0.288
SD-3 MM-DiT 7.7B 0.705 0.821 0.340
FLUX MM-DiT 16.9B 0.720 0.828 0.352
Table A1: Diptych Generation Comparisons. Quantitative comparisons of the diptych generation capabilities of various TTI models based on the total number of parameters, including the autoencoder, main network, and text encoder.
Our framework relies on the emerging property of the large-scale TTI model, FLUX, particularly its strong understanding of diptych property and the ability to represent diptych accurately. We verify this by synthesizing a total of $2100$ diptychs, using $20$ objects, each with a pair of two random prompts for each panel among $15$ prompts, and comparing the diptych generation performance with those of other previous TTI models. The prompt for diptych generation follows the setup mentioned in Sec. 3.1 of the main paper. We assessed the quality of each diptych by evaluating the interrelation and text alignment of each panel. This is measured through splitting the generated image in half and measuring DINO and CLIP-I scores between each panel, as well as the CLIP-T score between each panel and its description. The results are shown in Tab. A1, in which the diptych generation performance and total number of parameters including the autoencoder, main network, and text encoders are reported. These results exhibit the superior diptych generation capability of FLUX, where smaller models are insufficient. This allows us to extend to inpainting and propose a zero-shot subject-driven text-to-image generation method via diptych inpainting-based interpretation.
Appendix E Background Removal Ablation
We provide additional samples for the ablation study conducted with and without the background removal process $G_{\text{seg}}$ in Fig. A4. Consistent with the findings in the main paper, including the background leads to content leakage, where irrelevant elements such as background, pose, and location are mirrored in the generated results. This hinders the accurate reflection of the desired context described by the text and reduces diversity in pose and location. In contrast, removing the background and retaining only the subject information in the reference image on the left panel allows the generated outputs to better align with the desired context while exhibiting greater diversity in pose and location.”
Appendix F Reference Attention Enhancement Ablation
We further present the actual sample quality variations according to the reference attention rescaling factor $\lambda$ values to support the quantitative ablations in the main paper. These variations are visualized in Fig. A5. As seen in the qualitative results, the absence of reference attention enhancement ($\lambda=1.0$) can lead to a loss of fine details of the subject, resulting in subtle discrepancies such as the left eye of the backpack dog, the patch on its right eye, the fur color on the dog’s face, or the texture of the bear plushie’s fur. As the $\lambda$ value increases, these missed details are better preserved, leading to improved subject alignment performance. However, excessive enhancement can negatively impact the quality of the generated images, causing the subject to appear slightly blurred or exhibit minor color shifts.
Appendix G Stylized Image Generation
For stylized image generation, Diptych Prompting places the style image in the left panel and inpaints the right panel using the text prompt “A diptych with two side-by-side images of same style. On the left, {original image description}. On the right, replicate this style exactly but as {target image description} without attention enhancement ($\lambda=1.0$) for referencing only the stylistic elements except the content. Additional samples are provided in Fig. A6. Beyond the qualitative results, we also include quantitative comparisons using the same metrics (DINO, CLIP-T, CLIP-I) applied to a total of $2000$ generated images in Tab. A2. These images include $4$ samples per prompt and per style image, across $25$ prompts and $20$ style images collected from previous work [41]. As shown in the result, our method demonstrates comparable results to existing zero-shot style transfer methods specialized in stylized image generation, further proving the versatility of our approach.
Method DINO CLIP-I CLIP-T
RB-Mod [38] 0.295 0.598 0.372
IP-Adapter [50] 0.337 0.602 0.371
Diptych Prompting 0.357 0.623 0.349
Table A2: Stylized Image Generation Comparisons. Quantitative comparisons of stylized image generation with previous zero-shot methods.
Appendix H Subject-Driven Image Editing
Diptych Prompting is extended to the subject-driven image editing by placing the reference subject image on the left panel and the editing target image on the right panel in the incomplete diptych. By masking only the desired area in the right panel and applying diptych inpainting, the reference subject from the left panel is generated in the masked region on the right panel, resulting in the subject-driven image editing. Following the previous work [49], we conduct the subject-driven image editing with selected images from a subset of the MSCOCO [24] validation dataset, in which each image contains a bounding box and the bounding box is smaller than half of image size. We applied masking to the inside of the bounding box, enabling the generation of the reference subject within the specified region. More samples of various subjects and editing target images are available in Fig. A7.
Appendix I Limitations
Currently, FLUX is the only model with sufficient capability to effectively generate diptychs. However, as more advanced text-to-image (TTI) models become available, we anticipate that our method will be applicable to a wider range of models in the future. In line with advancements in other encoder-based zero-shot approaches, there is a need to explore multi-subject-driven text-to-image generation. We leave this exploration for future work. Furthermore, diptych generation requires the generated image to have an aspect ratio of $2:1$. Due to the limitation in the generatable resolution of FLUX, we were unable to produce the diptych image at a size of $2048\times 1024$ pixels and confirmed results up to $1536\times 768$ pixels, resulting in subject-driven image (right panel) being $768\times 768$ pixels in size. We expect that this issue can be easily addressed by utilizing super-resolution models such as ControlNet [51] or advanced TTI models for high-resolution image generation in the future.
Figure A2: Subject-Driven Text-to-Image Generation. More samples of subject-driven text-to-image generation using Diptych Prompting.
Figure A3: Subject-Driven Text-to-Image Generation. More samples of subject-driven text-to-image generation using Diptych Prompting..
Figure A4: $\bm{G_{\text{seg}}}$ Ablation. Qualitative comparisons with and without the background removal process.
Figure A5: $\bm{\lambda}$ Ablation. Qualitative transitions according to the varying $\lambda$ values. we control the $\lambda$ from $1.0$ (without reference attention enhancement) to $1.5$. For a detailed view, please zoom in.
Figure A6: Stylized Image Generation. More samples of stylized image generation using Diptych Prompting.
Figure A7: Subject-Driven Image Editing. More samples of subject-driven image editing using Diptych Prompting.

### v2

@@
 Our contributions can be summarized as follows:
-  • To the best of our knowledge, we are the first to approach zero-shot subject-driven text-to-image generation as an inpainting task, offering a new perspective by highlighting the diptych generation capabilities of FLUX.
+  • We propose a novel inpainting-based zero-shot subject-driven text-to-image generation approach without further training, offering a new perspective by highlighting the inherent diptych generation capabilities of FLUX.
   • We propose two techniques to prevent content leakage and reliably capture details in the target subject: isolating the subject from its background and enhancing attention weights between panels.
@@
 Moreover, optimizing on a small set of images may lead to overfitting on the new subject and catastrophic forgetting of prior knowledge which should be carefully prevented.
-The second group [50, 22, 47, 30, 46, 26, 31] addresses the limitations of the first by introducing image prompting, a framework that incorporates a reference image to guide the generated output.
-In this approach, a specialized image enco der [34] is utilized to support an additional image input alongside the text prompt, allowing the diffusion model to capture the characteristics of the reference subject while adhering to the context described by the text prompt.
-This approach enables zero-shot subject-driven text-to-image generation; however, it often lacks fidelity in the synthesized subject.
-We propose a novel inpainting-based zero-shot approach for subject-driven text-to-image generation using a large-scale TTI model with an inpainting module.
+The second group [55, 24, 51, 32, 50, 28, 33] addresses these limitations through image prompting, which utilizes a specialized image encoder [36] to incorporate a reference image alongside the text prompt to guide the generated output.
+Such methods enable zero-shot manner; yet, it often lacks target subject alignment.
+Other notable approaches [56, 17, 52] fine-tune the TTI model for joint-set image generation [2, 49, 46], and extend this to subject-driven generation or editing through inpainting.
+However, they still face training constraints, such as costs for dataset construction and training.
+Leveraging the inherent capability of a recent TTI model with an inpainting module, we propose a novel zero-shot inpainting-based approach without additional training.
 3 Method
@@
 We compare our method to previous zero-shot subject-driven text-to-image methods with encoder-based image prompting, including ELITE [51], BLIP-Diffusion [24], Kosmos-G [32], Subject-Diffusion [28], IP-Adapter [55], MS-Diffusion [50], and $\lambda$-Eclipse [33].
-The details of these models are provided in Appendix A.
+The details of these models are provided in appendix.
 Qualitative Results Our qualitative results are presented in Fig.
@@
 1, Diptych Prompting outperforms all baselines by a large margin ($p<0.01$ in the Wilcoxon signed-rank test), which is consistent with the qualitative results.
-Detailed information and full instructions about our human preference study are included in Appendix C.
+Detailed information and full instructions about our human preference study are included in appendix.
 Subject Align (%) Text Align (%)
@@
 IP-Adapter [55] $\dagger$ SD-XL 0.613 0.810 0.292
-IP-Adapter [50$\ddagger$ FLUX 0.561 0.725 0.351
-Diptych Prompting FLUX 0.688 0.758 0.345
+IP-Adapter [55$\ddagger$ FLUX 0.561 0.725 0.351
+Diptych Prompting FLUX 0.689 0.758 0.344
 Table 2: Quantitative Comparisons.
@@
 1.0 0.647 0.745 0.343
-1.3 0.688 0.758 0.345
+1.3 0.689 0.758 0.344
 1.5 0.670 0.750 0.342
@@
 We also assess the impact of varying the rescaling factor $\lambda$ on reference attention enhancement.
-Rescaling attention weights between the right panel query and the left panel key helps to capture fine details, thereby improving subject alignment.
+Rescaling attention weights between the right panel query and the left panel key helps to capture fine details, thereby improving subject alignment metrics.
 However, using too high values introduces excessive inductive bias, causing abnormal attention weights that negatively impact performance.
-Qualitative transitions with respect to $\lambda$ can be verified in Appendix F.
+Qualitative transitions with respect to $\lambda$ can be verified in appendix.
 4.4 Applications
@@
 Diptych Prompting successfully generates the stylistic image reflecting the style of the reference as shown in Fig.
-6, and the quantitative comparisons are available in Appendix G.
+6, and the quantitative comparisons are available in appendix.
 Subject-Driven Image Editing We further adapt our approach to support inpainting-based subject-driven image editing that modifies the target image with the specific subject.
@@
 This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) [No.
-2022R1A3B1077720], Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)], and the BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University in 2024.
-Appendix
-Appendix A Baselines
+2022R1A3B1077720], Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University), NO.
+RS-2022-II220959], the BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University in 2024, Samsung Electronics (IO221213-04119-01), and a grant from the Yang Young Foundation.
+Supplementary Material
+A Baselines
 We provide the details of the encoder-based image prompting baselines that we compared in human preference study, as well as in qualitative and quantitative evaluations.
@@
 In all experiments using IP-Adapter, regardless of the base model version, the conditioning scale is set to $0.6$.
-Figure A1: DreamBooth Comparisons. Quantitative comparisons to DreamBooth-LoRA with various rank values.
-Appendix B Subject-Driven Text-to-Image Generation
+Figure S1: DreamBooth Comparisons. Quantitative comparisons to DreamBooth-LoRA with various rank values.
+B Subject-Driven Text-to-Image Generation
 B.1 Evaluation Setting
@@
 The results are presented in Fig.
-A1, where our Diptych Prompting demonstrates superior performance across various model capacities.
+S1, where our Diptych Prompting demonstrates superior performance across various model capacities.
 B.3 Additional Results
 We include additional samples of Diptych Prompting in Fig.
-A2 and Fig.
-A3 for diverse objects and contexts.
+S2 and Fig.
+S3 for diverse objects and contexts.
 As demonstrated in the results, our methodology achieves high-quality image generation and satisfies both subject alignment and text alignment in a zero-shot manner by leveraging FLUX’s capabilities.
@@
 We also note that the fine details in the target subject are well reflected in the generated results, even for challenging subjects that previous zero-shot methods struggled with (e.g., robot toy, ‘Bon appetit’ bowl).
-Appendix C Human Preference Study
+C Human Preference Study
 Following the previous work [41], we perform the human preference study by pairwise comparison in two separate questionnaires for each aspect: subject alignment and text alignment.
@@
   • Which Machine-Generated Image is best described by the reference text?
-Appendix D Diptych Generation
+D Diptych Generation
 Model Arch Param DINO CLIP-I CLIP-T
@@
 FLUX MM-DiT 16.9B 0.720 0.828 0.352
-Table A1: Diptych Generation Comparisons.
+Table S1: Diptych Generation Comparisons.
 Quantitative comparisons of the diptych generation capabilities of various TTI models based on the total number of parameters, including the autoencoder, main network, and text encoder.
@@
 The results are shown in Tab.
-A1, in which the diptych generation performance and total number of parameters including the autoencoder, main network, and text encoders are reported.
+S1, in which the diptych generation performance and total number of parameters including the autoencoder, main network, and text encoders are reported.
 These results exhibit the superior diptych generation capability of FLUX, where smaller models are insufficient.
 This allows us to extend to inpainting and propose a zero-shot subject-driven text-to-image generation method via diptych inpainting-based interpretation.
-Appendix E Background Removal Ablation
+E Background Removal Ablation
 We provide additional samples for the ablation study conducted with and without the background removal process $G_{\text{seg}}$ in Fig.
-A4.
+S4.
 Consistent with the findings in the main paper, including the background leads to content leakage, where irrelevant elements such as background, pose, and location are mirrored in the generated results.
@@
 In contrast, removing the background and retaining only the subject information in the reference image on the left panel allows the generated outputs to better align with the desired context while exhibiting greater diversity in pose and location.”
-Appendix F Reference Attention Enhancement Ablation
+F Reference Attention Enhancement Ablation
 We further present the actual sample quality variations according to the reference attention rescaling factor $\lambda$ values to support the quantitative ablations in the main paper.
 These variations are visualized in Fig.
-A5.
+S5.
 As seen in the qualitative results, the absence of reference attention enhancement ($\lambda=1.0$) can lead to a loss of fine details of the subject, resulting in subtle discrepancies such as the left eye of the backpack dog, the patch on its right eye, the fur color on the dog’s face, or the texture of the bear plushie’s fur.
@@
 However, excessive enhancement can negatively impact the quality of the generated images, causing the subject to appear slightly blurred or exhibit minor color shifts.
-Appendix G Stylized Image Generation
+G Stylized Image Generation
 For stylized image generation, Diptych Prompting places the style image in the left panel and inpaints the right panel using the text prompt “A diptych with two side-by-side images of same style.
@@
 Additional samples are provided in Fig.
-A6.
+S6.
 Beyond the qualitative results, we also include quantitative comparisons using the same metrics (DINO, CLIP-T, CLIP-I) applied to a total of $2000$ generated images in Tab.
-A2.
+S2.
 These images include $4$ samples per prompt and per style image, across $25$ prompts and $20$ style images collected from previous work [43].
@@
 Diptych Prompting 0.357 0.623 0.349
-Table A2: Stylized Image Generation Comparisons. Quantitative comparisons of stylized image generation with previous zero-shot methods.
-Appendix H Subject-Driven Image Editing
+Table S2: Stylized Image Generation Comparisons. Quantitative comparisons of stylized image generation with previous zero-shot methods.
+H Subject-Driven Image Editing
 Diptych Prompting is extended to the subject-driven image editing by placing the reference subject image on the left panel and the editing target image on the right panel in the incomplete diptych.
@@
 More samples of various subjects and editing target images are available in Fig.
-A7.
-Appendix I Limitations
+S7.
+I Limitations
 Currently, FLUX is the only model with sufficient capability to effectively generate diptychs.
@@
 We expect that this issue can be easily addressed by utilizing super-resolution models such as ControlNet [57] or advanced TTI models for high-resolution image generation in the future.
-Figure A2: Subject-Driven Text-to-Image Generation. More samples of subject-driven text-to-image generation using Diptych Prompting.
-Figure A3: Subject-Driven Text-to-Image Generation. More samples of subject-driven text-to-image generation using Diptych Prompting..
-Figure A4: $\bm{G_{\text{seg}}}$ Ablation. Qualitative comparisons with and without the background removal process.
-Figure A5: $\bm{\lambda}$ Ablation.
+Figure S2: Subject-Driven Text-to-Image Generation. More samples of subject-driven text-to-image generation using Diptych Prompting.
+Figure S3: Subject-Driven Text-to-Image Generation. More samples of subject-driven text-to-image generation using Diptych Prompting..
+Figure S4: $\bm{G_{\text{seg}}}$ Ablation. Qualitative comparisons with and without the background removal process.
+Figure S5: $\bm{\lambda}$ Ablation.
 Qualitative transitions according to the varying $\lambda$ values.
@@
 For a detailed view, please zoom in.
-Figure A6: Stylized Image Generation. More samples of stylized image generation using Diptych Prompting.
-Figure A7: Subject-Driven Image Editing. More samples of subject-driven image editing using Diptych Prompting.
+Figure S6: Stylized Image Generation. More samples of stylized image generation using Diptych Prompting.
+Figure S7: Subject-Driven Image Editing. More samples of subject-driven image editing using Diptych Prompting.